StonyBrook Institute for Mathematical Sciences

SUNY, Stony Brook, New York 11794-3651 telephone: (631) 632-7318 email: IMS@math.sunysb.edu

PREPRINTS IN THIS SERIES:

1998/1. S. Zakeri, Biaccessiblility in Quadratic Julia Sets I: The Locally-Connected Case; D. Schleicher & S. Zakeri, On Biaccessible Points in the Julia Set of a Cremer Quadratic Polynomial; S. Zakeri, Biaccessiblility in Quadratic Julia Sets II: The Siegel and Cremer Cases.

/2. V. Kaimanovich, The Poisson Formula for Groups with Hyperbolic Properties. /3. E. Bedford & M. Jonsson, Regular Polynomial Endomorphisms of C^{k}. /4. S. Zakeri, On Dynamics of Cubic Siegel Polynomials. /5. M. Yampolsky, The Attractor of Renormalization and Rigidity of Towers of Critical Circle Maps. /6. C. Bishop, Non-removable Sets for Quasiconformal and Locally BiLipschitz Mappings in R^{3}. /7.*J. J. P. Veerman, M. M. Peixoto, A. C. Rocha, & S. Sutherland, On Brillouin Zones. Comm.

Math. Phys. 212 (2000), 725-744. /8. M. Yampolsky & S. Zakeri, Mating Siegel Quadratic Polynomials. /9. J. Milnor & C. Tresser, On Entropy and Monotonicity for Real Cubic Maps (with an appendix by

A. Douady & P. Sentenac). Revised June 1999. /10. S. Zakeri, Dynamics of Singular Holomorphic Foliations on the Complex Projective Plane. /11. J. Kahn, Holomorphic Removability of Julia Sets. /12. D. Schleicher, On Fibers and Local Connectivity of Compact Sets in C. /13. D. Schleicher, I: On Fibers and Local Connectivity of Mandelbrot and Multibrot Sets.

II: On Fibers and Renormalization of Julia Sets and Multibrot Sets.

1999/1. A. Epstein, Infinitesimal Thurston Rigidity and the Fatou-Shishikura Inequality.

/2. V. Kaloshin, Generic Diffeomorphisms with Superexponential Growth of Number of Periodic Orbits. /3. J. Milnor, Periodic Orbits, Externals Rays and the Mandelbrot Set: An Expository Account. /4. X. Buff & C. Henriksen, Scaling Ratios and Triangles in Siegel Disks. /5. K. Pilgrim, Dessins d’enfants and Hubbard Trees. /6. W. de Melo & A. A. Pinto, Rigidity of C^{2} Infinitely Renormalizable Unimodal Maps. /7. N. Shah & B. Weiss, On Actions of Epimorphic Subgroups on Homogeneous Spaces;

B. Weiss, Unique Ergodicity on Compact Homogeneous Spaces. /8. E. Lindenstrauss & B. Weiss, On Sets Invariant under the Action of the Diagonal Group. /9. D. Schleicher & J. Zimmer, Dynamic Rays for Exponential Maps. /10. J. J. P. Veerman & B. D. Stošić, On the Dimensions of Certain Incommensurably Constructed Sets.

2000/1. J. Hubbard & P. Papadopol, Newton’s Method Applied to Two Quadratic Equations in C^{2} Viewed as

a Global Dynamical System. /2. A. de Carvalho & T. Hall, Pruning, Kneading and Thurston’s Classification of Surface Homeomor-

phisms. /3. T. Lundh, In Search of an Evolutionary Coding Style. /4. D. Schleicher, Attracting Dynamics of Exponential Maps. /5. F. Loray & J. Rebelo, Stably Chaotic Rational Vector Fields on CP^{n}. /6. C. L. Peterson & s. Zakeri, On the Julia Set of a Typical Quadratic Polynomial with a Siegel Disk. /7. Y. Minsky, Bounded Geometry for Kleinian Groups. /8. Y.Minsky & B. Weiss, Nondivergence of Horocyclic Flows on Moduli Space. /9. J. Rivera-Letelier, Rational Maps with Decay of Geometry: Rigidity, Thurston’s Algorithm and Local

Connectivity. /10. A. Avila, M. Martens & W. de Melo, On the Dynamics of the Renormalization operator.

Continued on back cover.

*Starred papers have already appeared in the journal cited.

StonyBrook Institute for Mathematical Sciences

SUNY, Stony Brook, New York 11794-3651 telephone: (631) 632-7318 email: IMS@math.sunysb.edu

PREPRINTS IN THIS SERIES: Continued from inside cover.

/11. F. Ferreira & A. A. Pinto, Explosion of Smoothness From a Point to Everywhere for Conjugacies

Between Diffeomorphisms on Surfaces. 2001/1. E. de Faria, W. de Melo & A. Pinto, Global Hyperbolicity of Renormalization for C^{r} Unimodal

Mappings. /2. A. de Carvalho & T. Hall, The Forcing Relation for Horshoe Braid Types. /3. G. Birkhoff, M. Martens & C. Tresser, On the Scaling Structure of Period Doubling. /4. A. Epstein & M. Yampolsky, A Universal Parabolic Map. /5. V. Kaimanovich & M. Lyubich, Conformal and Harmonic Measures on Laminations Associated with

Rational Maps. /6. H. Miyachi, Cusps in Complex Boundaries of One-Dimensional Teichmuller Space. /7. F. P. Gardiner, J. Hu & N. Lakic, Earthquake Curves. /8. Jun Hu, Earthquake Measure and Cross-ratio Distortion. /9. J. C. Rebelo & R. R. Silva, The multiple ergodicity of non-discrete subgroups of Diff^{w}(S^{1}). /10. A. de Carvalho & T. Hall, How to prune a horshoe. /11. G. Tomanov & B. Weiss, Closed orbits for actions of maximal tori on homogeneous spaces. /12. J. Rivera-Letelier, Espace hyperbolique p-adique et dynamique des fonctions rationnelles. /13. M. Lenci, Billiards with an infinite cusp. /14. J. T. Tyson, On the conformal dimensions of quasiconvex post-critically finite self similar sets. /15. A. Avila, M. Lyubich & W. de Melo, Regular or stochastic dynamics in real analytic families of

unimodal maps. 2002/1. A. de Carvalho & T. Hall, Braid forcing and star-shaped train tracks.

/2. S. Zakeri, External rays and the real slice of the mandelbrot set. /3. J. C. Rebelo, Complete polynomial vector fields on C^{2}, Part I. /4. A. de Carvalho & M. Paternain, Monotone quotients of surface diffeomorphisms. /5. S. Zakeri, David maps and Hausdorff Dimension.

2003/1. A. Carocca & R. E. Rodríguez, Jacobians with group actions and rational idempotents. /2. S. R. Simanca, Heat Flows for Extremal Kähler Metrics.

The IMS preprint series began in 1990. Publications in the series have been posted and made available on the World Wide

Web at http://www.math.sunysb.edu/preprints.html, as well as via email or anonymous ftp from ftp.math.sunysb.edu.

Send an email message containing the single line send info to preprints@math.sunysb.edu for further details.

This preprint series is supported in part by the National Science Foundation